# Quantum Simulation of the Unruh Temperature Via the Thermal Properties of Virtually Evolving Bose-Einstein Condensates

### v2

@@ II.4 Determination of the New Coupling Frequency $g_{ch}$
Before computing the eigenspectrum of (3), we need to determine the formula for the new coupling frequency $g_{ch}$. To do this, we utilize the expression for the coupling frequency derived by Hu et al. by exploring the mathematical correspondence between the transformation resulting from the analytical solution of the time evolution under the Hamiltonian in (2) and the Rindler transformation. For a final number of emitted matter-waves, this coupling frequency takes the form:
@@
    The final number of emitted matter waves is proportional to a characteristic average thermal wavelength ($\lambda_{ch}$), which is a function of the characteristic acceleration ($A_{ch}$) characterizing the emitted matter waves simulating the Unruh temperature: $\lambda_{ch}\propto A_{ch}^{-1/2}$.
@@
Based on the first assumption, the characteristic can be expressed in terms of the characteristic acceleration, thus determining the characteristic angular frequency ($\omega_{ch}$). Using :
@@
The final time ($\tau_{ch}$) is then deduced using the saturated Heisenberg uncertainty principle, $\Delta E_{ch}\Delta\tau={\hbar}/2$, where the energy of the characteristic emitted waves is $E_{ch}=n\hbar\omega_{ch}/2$:

### v3

@@ II.4 Determination of the New Coupling Frequency $g_{ch}$
Before computing the eigenspectrum of (3), we need to determine the formula for the new coupling frequency $g_{ch}$. To do this, we utilize the expression for the coupling frequency derived by Hu et al. by exploring the mathematical correspondence between the transformation resulting from the analytical solution of the time evolution under the Hamiltonian in (2) and the Rindler transformation. For a characteristic number of emitted matter-waves, this coupling frequency takes the form:
@@
    The characteristic number of emitted matter waves is proportional to a characteristic average thermal wavelength ($\lambda_{ch}$), which is a function of the characteristic acceleration ($A_{ch}$) characterizing the emitted matter waves simulating the Unruh temperature: $\lambda_{ch}\propto A_{ch}^{-1/2}$.
@@
Based on the first assumption, the characteristic number of emitted matter waves can be expressed in terms of the characteristic acceleration, thus determining the characteristic angular frequency ($\omega_{ch}$). Using :
@@
The characteristic time ($\tau_{ch}$) is then deduced using the saturated Heisenberg uncertainty principle, $\Delta E_{ch}\Delta\tau={\hbar}/2$, where the energy of the characteristic emitted matter waves is $E_{ch}=n\hbar\omega_{ch}/2$: